# A Benchmark of Causal vs. Correlation AI for Predictive Maintenance

### v1

@@ 3.2.3. Probabilistic and Causal Approaches
L4: Bayesian Logistic Regression Probabilistic logistic model Bayesian inference with uncertainty quantification
L5: Causal Model Causal structural model Causal graph–based modeling
L6: Rule-Based Causal Model Hand-crafted causal rules Learned thresholds on mechanistic rules
L7: PyMC BART Model Bayesian tree ensemble Non-parametric sum-of-trees with causal features
@@ 3.2.4. Validation Strategy
To ensure results are robust and not artifacts of a single train-test split, we implemented a comprehensive validation strategy. All models (L0 through L5) were trained and evaluated across five different random seeds (42, 43, 44, 45, and 46). For each seed, the dataset is randomly shuffled before the 80-20 stratified split, producing different training and test sets while maintaining the 3.3% failure rate in both partitions.
@@
Additionally, we examine the generalization gap; the difference between training-set performance and test-set performance. A large gap indicates overfitting: the model has memorized training data patterns that do not generalize. We report both training and test savings percentages, with the gap measured in percentage points.
@@ 4. Results and Analysis
Table 6 summarizes test-set performance for all six models, averaged across five random seeds. Each model was evaluated on 2,000 held-out machines containing 66 deterministic failures, with a baseline reactive maintenance cost of 1,650,000 USD.
@@
L1: Balanced LogReg 743,000 907,000 55.0% 81.8% 13.5% 0.232 54 346 12
L2: Cost-Aware LogReg 794,000 856,000 51.9% 83.3% 10.1% 0.181 55 488 11
L3: Cost-Aware Tree 572,500 1,077,500 65.3% 87.9% 26.0% 0.401 58 165 8
L4: Bayesian LogReg 878,000 772,000 53.2% 81.8% 13.5% 0.232 54 346 12
L5: Causal 492,500 1,157,500 70.2% 87.9% 92.1% 0.899 58 5 8
L6: Rule-Based Causal 477,000 1,173,000 71.1% 100.0% 18.3% 0.310 66 294 0
L7: PyMC BART 541,000 1,109,000 67.2% 89.4% 29.4% 0.442 59 142 7
Table 6. Model performance, costs, and savings
@@
  • Models L1, L2, and L4—all variations of logistic regression—cluster around a 51-55% cost reduction, saving between 772,000 USD and 907,000 USD annually.
  • Model L3, leveraging non-linear decision trees, achieves a 65.3% cost reduction with 1,077,500 USD in annual savings.
  • Model L5, using formal causal inference, achieves a 70.2% cost reduction with 1,157,500 USD in annual savings while keeping false alarms in the single digits, close to zero.
  • Model L6’s rule-based thresholds push recall to 100% and edge out L5 in total cost, but do so by triggering 294 inspections—nearly five times more than the cost-aware tree—which would overwhelm maintenance crews.
  • Model L7’s PyMC BART sits between L3 and L5: it matches their recall and yet generates more false positives because it lacks hard structural constraints, illustrating the limits of “feature generation only” strategies.
@@ 4.1. Detailed Model Analysis
L0: No-Skill Baseline. As expected, the no-skill model does not contribute anything. By always predicting “no failure,” it achieves the naive accuracy of 96.7% (equal to the base rate of non-failures) while catching zero actual failures. All 66 failures in the test set occur unexpectedly, each incurring the full 25,000 USD emergency cost, totaling 1,650,000 USD. This model serves only as the reference point against which all other approaches must be measured.
L1: Balanced Logistic Regression. The first predictive model demonstrates that even simple linear models can extract meaningful signal from sensor data. Model L1 catches 54 out of 66 failures (81.8% recall), preventing 1,350,000 USD in potential downtime costs. These 54 true positives cost 5,000 USD each in preventive maintenance, totaling 270,000 USD.
However, the model also generates 346 false positives—unnecessary maintenance interventions on healthy machines. At 500 USD each, these false alarms cost 173,000 USD in wasted labor. The 12 missed failures (false negatives) each incur the full 25,000 USD emergency cost, totaling 300,000 USD. Summing all components yields total cost of 743,000 USD, representing 907,000 USD in savings compared to reactive maintenance—a 55% cost reduction.
The model’s 13.5% precision indicates that only about 1 in 7 alarms corresponds to a real failure. While this may seem poor, it is acceptable given the cost structure: each real failure costs 25,000 USD while each false alarm costs only 500 USD. The model can afford 50 false alarms per true positive and still break even economically.
L2: Cost-Aware Logistic Regression. Model L2 applies the same logistic regression architecture as L1 but optimizes the decision threshold explicitly for business costs. The threshold optimization pushes the decision boundary lower, making the model more aggressive in predicting failures. This increases recall slightly to 83.3% (55 out of 66 failures caught) but generates substantially more false alarms—488 instead of 346.
The cost breakdown reveals why this strategy backfires. The 55 true positives cost 275,000 USD, the 488 false positives cost 244,000 USD, and the 11 false negatives cost 275,000 USD, totaling 794,000 USD. While the model catches one additional failure compared to L1, it generates 142 additional false alarms, adding 71,000 USD in unnecessary intervention costs. Net savings of 856,000 USD are actually 51,000 USD lower than L1.
This counterintuitive result illustrates a subtle point: threshold optimization on the training set can lead to overfitting of the decision boundary. The optimal threshold for training data may be too aggressive for test data if the model’s probability calibration is poor [11]. Model L2’s probabilities are not well-calibrated—many predicted “high risk” machines do not actually fail—so lowering the threshold excessively inflates false alarms.
L3: Cost-Aware Decision Tree. Model L3 represents the best correlation-based approach by combining non-linear modeling with constrained complexity. The decision tree architecture naturally captures feature interactions such as “heat dissipation failure occurs when temperature differential is low AND rotational speed is low.” Regularization hyperparameters (maximum depth 5, minimum 50 samples per split, minimum 20 samples per leaf) prevent the tree from memorizing training data noise.
This model catches 58 out of 66 failures (87.9% recall), a substantial improvement over L1’s 54. Critically, it does so with only 165 false alarms—less than half of L1’s 346 false positives. This superior precision (26.0%, meaning about 1 in 4 alarms is real) reflects the model’s ability to identify genuine failure patterns rather than overfitting to spurious correlations.
The cost breakdown is favorable: 58 true positives at 5,000 USD each cost 290,000 USD, 165 false positives at 500 USD each cost 82,500 USD, and 8 missed failures at 25,000 USD each cost 200,000 USD. Total cost of 572,500 USD yields 1,077,500 USD in annual savings—a 65.3% cost reduction. This represents the ceiling of what correlation-based machine learning can achieve on this problem.
L4: Bayesian Logistic Regression. Model L4 applies Bayesian inference to quantify prediction uncertainty, but this additional sophistication does not translate to improved test-set performance. The model achieves 81.8% recall and 13.5% precision, identical to L1’s results. Total cost of 878,000 USD yields 772,000 USD in savings—better than reactive maintenance but worse than simple balanced logistic regression (L1).
The modest performance suggests that parameter uncertainty, while theoretically valuable, does not substantially improve predictions when the underlying model class (linear logistic regression) has limited expressiveness. The Bayesian framework provides interpretability—we can examine posterior distributions to assess which features have reliably non-zero effects—but does not overcome the fundamental limitation that linear models cannot capture the complex non-linear interactions present in failure mechanisms.
The computational cost of Bayesian inference (several minutes for MCMC sampling versus seconds for standard logistic regression) further reduces practical appeal unless the uncertainty quantification itself is valuable for decision-making. For this predictive maintenance application, actionable predictions matter more than prediction intervals.
L5: Causal Model. Model L5 achieves the highest performance across all metrics, demonstrating the value of causal reasoning. The model catches 58 out of 66 failures (87.9% recall), matching Model L3’s recall. However, it does so with only 5 false positives compared to L3’s 165—a dramatic 97% reduction in false alarms. Precision of 92.1% means that more than 9 out of 10 alarms correspond to real failures.
This exceptional precision-recall balance translates to substantial cost savings. The 58 true positives cost 290,000 USD, the mere 5 false positives cost only 2,500 USD (negligible), and the 8 missed failures cost 200,000 USD. Total cost of 492,500 USD yields 1,157,500 USD in annual savings—a 70.2% cost reduction and the highest financial value among all models.
The model’s superiority stems from its causal architecture. By encoding domain knowledge in the causal graph and deriving physically meaningful features (temperature differential for thermal stress, power for energy dissipation, overstrain for mechanical stress), the model learns relationships that generalize robustly. When the model predicts a failure, it does so because a specific causal mechanism has been activated—not because of a spurious correlation in the training data.
For example, a prediction might be traced to high overstrain: tool wear has reached 215 minutes while torque is 56 Nm, producing an overstrain index of 12,040 that exceeds the safe threshold of 11,000. This level of interpretability enables targeted interventions: the maintenance team can replace the tool before reaching the dangerous threshold, addressing the root cause directly.
L6: Rule-Based Causal Model. The threshold-based rule system provides a useful upper bound on recall. By declaring a failure whenever any mechanism is even close to its safety margin, the rules catch all 66 deterministic failures (100% recall) and reduce total cost to 477,000 USD. However, the policy produces 294 false alarms—orders of magnitude more than the five false alarms produced by L5. In practice that would translate into hundreds of unnecessary stoppages per week. We therefore treat L6 as a “safety-first” bound that validates the physics of the rules, not as a candidate deployment policy.
L7: PyMC BART Model. BART delivers 1,109,000 USD in savings (67.2%) with recall of 89.4%, but its precision drops to 29.4% (142 false positives). This outcome answers potential questions around automatic feature generation: even a powerful Bayesian ensemble that can discover complex interactions still benefits from graph-regularized features. Without explicit causal constraints the model develops conservative decision boundaries, trading cost savings for a larger inspection queue.
4.1.1. Comparative Analysis: Causal vs. Correlation
The performance gap between Model L5 (causal) and Model L3 (best correlation) quantifies the value of causal reasoning. Model L5 saves an additional 80,000 USD annually (7.4% improvement in savings percentage) while generating 160 fewer false alarms—a 97% reduction in operational disruption. L6 shows that we can squeeze a few more dollars out of the problem by firing every possible alarm, but at the cost of 294 inspections; L5 reaches nearly the same savings with almost no false positives.
This improvement cannot be attributed to model complexity. L3 is a single, carefully regularized decision tree rather than a gradient-boosted ensemble, whereas L5 couples do-calculus adjustments with gradient boosting and L7 adds a Bayesian additive regression tree. When we train powerful ensembles (gradient boosting and BART) on raw features without causal guidance, they rediscover similar splits yet inflate the false-positive rate. The difference lies in the features themselves. Model L3 learns from raw sensor readings, discovering correlations like “high torque readings are associated with failure.” Model L5 learns from causally derived features, understanding mechanisms like “high torque combined with worn tools causes mechanical overstrain failure,” while Model L7 demonstrates that even a flexible Bayesian ensemble still benefits from causal priors.
When operating conditions shift slightly—perhaps a new batch of machines has different temperature baselines, or a process change alters typical torque ranges—Model L3’s correlations may degrade, while Model L5’s causal relationships remain valid. The physics of mechanical overstrain do not change when ambient temperature drifts, whereas the correlation between specific torque values and failure does.
This distinction has profound implications for model deployment and longevity. Model L3 requires periodic retraining as data distributions shift. Model L5’s causal structure provides inherent robustness to distribution changes, as causal relationships are invariant properties of the system rather than contingent patterns in historical data.
@@ 4.1.2. Generalization and Overfitting Assessment
To assess whether observed performance reflects genuine model capability or lucky train-test splits, we examine the generalization gap—the difference between training performance and test performance. Table 7 shows training savings percentage, test savings percentage, and the gap in percentage points for predictive models.
@@
L1: Balanced LogReg 56.9% 55.0% 1.97
L2: Cost-Aware LogReg 55.8% 51.9% 3.93
L3: Cost-Aware Tree 69.4% 65.3% 4.11
L4: Bayesian LogReg 56.9% 53.2% 3.70
L5: Causal 72.8% 70.2% 2.60
L6: Rule-Based Causal 73.5% 71.1% 2.40
L7: PyMC BART 69.1% 67.2% 1.87
Table 7. Train–test savings comparison. Gap represents the difference between training and test savings percentages; smaller gaps indicate better generalization.
All models exhibit modest generalization gaps below 5 percentage points, which we set as an acceptable bound because plant operators typically budget that amount of degradation when deploying predictive maintenance models across production lines [8]. The multi-seed experiment (five independent splits) confirmed that variance stays within this tolerance without any additional regularization. Model L3 shows the largest gap at 4.11 percentage points, suggesting slight over-optimization to training data, yet its test performance remains the highest among correlation-based approaches.
Logistic regression (L1) records the smallest gap at 1.97 percentage points, with BART close behind at 1.87 percentage points. Model L5’s 2.60 point gap is only marginally larger while delivering far superior cost metrics, reinforcing the hypothesis that causal features generalize more robustly than raw correlations even if they do not literally yield the lowest generalization gap.
These findings are reinforced by multi-seed validation. Across five different random splits, model rankings remained completely stable: L5 outperformed L3, which outperformed L1, which outperformed L2, which outperformed L0. This consistency demonstrates that observed performance differences reflect true model capabilities rather than chance variations in data partitioning.
@@ 5. Conclusion
This study evaluated six predictive maintenance models on a CNC machine failure dataset, comparing traditional correlation-based machine learning approaches with formal causal inference methods. Results indicate that the causal AI approach (Model L5) achieved superior performance on business-relevant metrics, delivering 1,157,500 USD in annual cost savings (70.2% reduction) while maintaining high recall (87.9%) and exceptional precision (92.1%). The causal model outperformed the best correlation-based decision tree (Model L3) by 80,000 USD annually while reducing false alarms by 97%.
@@
    Causal superiority: When domain knowledge enables construction of causal graphs, formal causal inference methods can provide both superior predictive performance and interpretable root-cause explanations for predictions.
@@
    Bayesian ensembles still benefit from structure: The PyMC BART model validates that automatic feature generation can approach causal performance, but without graph-informed constraints, it issues substantially more false positives.
@@
    Robust generalization: Causally-derived features demonstrate stronger generalization (2.6 percentage point train-test gap) compared to raw sensor correlations (4.1 percentage points for the best correlation model), consistent with theoretical expectations that causal relationships are invariant under distribution shifts.
@@
    Operational benefits: The dramatic reduction in false alarms (97% fewer than correlation-based approaches) minimizes operational disruption while maintaining safety through high recall.
@@
    Static analysis: This study evaluates models on historical data without real-world deployment. Practical implementation would require monitoring for distribution drift, model degradation, and integration with existing maintenance management systems.
@@
Future research directions include: (1) validation across multiple manufacturing domains and equipment types; (2) investigation of causal discovery algorithms to automate DAG construction; (3) development of online learning approaches for continual model adaptation; (4) integration with IoT sensor networks and SCADA systems for real-time deployment; and (5) extension to multi-component systems with cascading failure modes.

### v2

@@ 3.2.4. Causal Model
L4: Random Forest Decision tree ensemble Average the predictions of multiple trees
L5: AdaBoost Boosted weak learner ensemble Sequentially reweight misclassified samples
L6: LightGBM Gradient boosting machine Histogram-based leaf-wise tree growth with boosting
L7: Bayesian SCM Bayesian structural causal model Explicit failure mechanisms with Noisy-OR combination
@@ 3.2.5. Validation Strategy
To ensure results are robust and not artifacts of a single train-test split, we implemented a comprehensive validation strategy. All models (L0 through L7) were trained and evaluated across five different random seeds (42, 43, 44, 45, and 46). For each seed, the dataset is randomly shuffled before the 80-20 stratified split, producing different training and test sets while maintaining the 3.3% failure rate in both partitions.
@@
Additionally, we examine the generalization gap; the difference between training-set performance and test-set performance. A large gap indicates overfitting: the model has memorized training data patterns that do not generalize. We report both training and test savings percentages, with the gap measured in percentage points. A test performance within 5 percentage points of training performance we deem acceptable (following [8]).
@@ 4. Results and Analysis
Table 6 summarizes test-set performance for all models, averaged across five random seeds. Each model was evaluated on 2,000 held-out machines containing 66 deterministic failures, with a baseline reactive maintenance cost of 1,650,000 USD.
@@
L1: Balanced LogReg 796,600 853,400 51.7% 77.3% 13.3% 0.227 51 333 15
L2: Cost-Aware LogReg 813,000 837,000 50.7% 82.4% 9.8% 0.175 54 502 11
L3: Cost-Aware Tree 576,900 1,073,100 65.0% 90.0% 23.6% 0.364 59 229 6
L4: Random Forest 481,500 1,168,500 70.8% 96.4% 23.6% 0.379 64 207 2
L5: AdaBoost 531,500 1,118,500 67.8% 95.2% 18.9% 0.314 63 275 3
L6: LightGBM 510,900 1,139,100 69.0% 91.5% 31.2% 0.364 60 138 6
L7: Bayesian SCM 554,900 1,095,100 66.4% 90.3% 26.2% 0.395 59 193 6
Table 6. Model performance, costs, and savings, averaged over validation splits. TP, FP, and FN counts are rounded to the nearest integer.
@@
  • Models L1 and L2—variations of logistic regression—cluster around a 50–52% cost reduction, saving between 837,000 USD and 853,400 USD annually.
  • Model L3, leveraging non-linear decision trees, achieves a 65.0% cost reduction with 1,073,100 USD in annual savings.
  • Model L4 (Random Forest) achieves the best performance among all models with a 70.8% cost reduction and 1,168,500 USD in annual savings and the highest recall (96.4%), catching 64 out of 66 failures with only 2 missed.
  • Model L5 (AdaBoost) achieves a 67.8% cost reduction with 1,118,500 USD in annual savings and 95.2% recall (63 of 66 failures caught).
  • Model L6 (LightGBM) achieves a 69.0% cost reduction with 1,139,100 USD in annual savings. Though the model predicted considerably fewer false positives compared to random forest, its lower recall still results in slightly higher costs due to missed failures.
  • Model L7 (Bayesian SCM) achieves a 66.4% cost reduction with 1,095,100 USD in annual savings and 90.3% recall. While this places it behind the best ensemble models on raw cost metrics, it is the only model capable of attributing predicted failures to specific types, as discussed in Section 4.1.1.
@@ 4.1. Detailed Model Analysis
The no-skill baseline (L0) catches zero failures and incurs the full 1,650,000 USD reactive maintenance cost. This serves as baseline for evaluating other models. Logistic regression models and their variants (L1, L2) resulted in 47–55% cost reduction relative to the baseline. Their suboptimal results are not surprising considering the nonlinear true data generating process and the linear decision boundary these models use to make predictions.
The non-linear tree-based correlation-based models (L3, L4, L5, L6) perform substantially better with 65–71% cost reduction. Random Forest (L4) had the best performance as it misses only 2 failures with a reasonable number of false positives. It is worth noting that the underlying data generating process of this dataset is itself a decision tree. Yet ensemble models, even equipped with a weak learner like AdaBoost, outperform a single decision tree. This demonstrates the strength of ensemble methods as they aggregate multiple estimators and can overcome the single tree’s tendency to overfit. On the other hand, a single decision tree has better interpretability since one can trace the decision path down the tree to infer the failure mode. This property is lost in ensemble models, in exchange for superior accuracy and variance reduction.
The causal model, Bayesian SCM (L7), achieves a 66.4% cost reduction and the highest recall among non-ensemble models (90.3%). Compared to the best of the correlation-based models, the causal model suffers a modest performance loss in exchange for stronger interpretability. That is, when the model predicts a failure, it simultaneously attributes that failure to the most likely causes. This allows the maintenance team to address the specific failure mechanism identified by the model, which can be a favorable trade-off in real-world maintenance settings where understanding why a machine will fail is as important as knowing that it will.
4.1.1. Failure Attribution
Among the correlation-based models, the decision tree offers the strongest level of interpretability. One may trace the decision path, record the features that drive the classification and infer the failure types. This, however, requires additional effort to perform the backtracing. Other correlation-based models like ensemble methods present a deeper challenge, as the reasoning is masked behind the aggregation of many weak learners. The causal model, with the domain knowledge encoded in the structure equations, has the inherent ability to attribute failure without needing interpretation efforts.
We examine the quality of failure attribution reported by the Bayesian SCM. Recall that the model was trained using only binary machine failure labels without knowing specific failure types (TWF, HDF, PWF, or OSF). Inside the model it uses structural equations to estimate the probability of each failure mode. When the model predicts a failure, we attribute the failure to the type with the greatest probability. We evaluate the attribution results against the ground truth failure type labels available in the dataset. Table 7 reports the number of ground truth failures in the test set (seed 42), the number detected by the model, and the number correctly attributed among those detected.
Failure Type Ground Truth Detected Correctly Attributed
TWF 7 3 1
HDF 25 25 25
PWF 15 15 15
OSF 22 22 22
Table 7. Per-type failure attribution performance on the test set (seed 42). Attribution analysis is reported for a single representative seed as failure type labels were not recorded in the averaged results.
For those detected failures, the Bayesian SCM achieves 100% correct attribution for HDF, PWF, and OSF types. This encouraging result shows that the structural equations successfully captured the physical mechanisms of each failure type, especially considering that the model was never given the failure type labels during training. TWF attribution, however, presents a unique challenge rooted in the data-generating process itself. As described in the dataset documentation [4], TWF involves two sequential random checks. Firstly, each machine is assigned a random wear threshold between 200–240 minutes. Secondly, upon crossing that threshold, whether the machine actually fails or receives a tool replacement is determined by another independent random assignment (46 failures out of 120 threshold crossings in the entire dataset). The first random check can in principle be approximated by a probabilistic model. However, the second check is entirely independent of all sensor readings. This introduces irreducible noise which cannot be resolved by any deterministic classification model. As a result, the model achieves lower performance on TWF attribution compared to other failure types. This is expected given the stochasticity in the data-generating process. For the remaining three failure types, which are governed by deterministic threshold rules on observable sensor features, the Bayesian SCM provides accurate attribution.
@@ 4.1.2. Generalization and Overfitting Assessment
We performed the multi-seed validation tactic as explained in 3.2.5. Table 8 shows training savings percentage, test savings percentage, and the gap in percentage points for predictive models.
@@
L1: Balanced LogReg 57.2% 51.7% 5.44
L2: Cost-Aware LogReg 56.0% 50.7% 5.26
L3: Cost-Aware Tree 69.4% 65.0% 4.36
L4: Random Forest 73.3% 70.8% 2.44
L5: AdaBoost 69.6% 67.8% 1.82
L6: LightGBM 72.6% 69.0% 3.60
L7: Bayesian SCM 69.5% 66.4% 3.17
Table 8. Train–test savings comparison averaged over validation splits. Gap represents the difference between training and test savings percentages; smaller gaps indicate better generalization.
Most models exhibit modest generalization gaps below 5 percentage points, which we set as an acceptable bound because plant operators typically budget that amount of degradation when deploying predictive maintenance models across production lines [8]. Notably, logistic regression variants (L1, L2) are exceptions which marginally exceed the acceptable bound at 5.44 and 5.26 percentage points respectively. This failure to generalize is not surprising given the inconsistency between their linear decision boundary and the nonlinear ground truth. The remaining models all fall within the acceptable range. AdaBoost (L5) records the smallest gap at 1.82 percentage points, and Random Forest (L4) follows at 2.44 points. The Bayesian SCM (L7) achieves a 3.17 point gap, consistent with the expectation that causal relationships are less susceptible to distributional idiosyncrasies in the training sample. The multi-seed validation confirmed stable model rankings across all five splits, demonstrating that observed performance differences reflect genuine model capabilities rather than chance variation in data partitioning.
@@ 5. Conclusion
This study evaluated eight predictive maintenance models on a CNC machine failure dataset, comparing traditional correlation-based machine learning approaches with Bayesian structural causal methods. Among correlation-based models, ensemble methods consistently achieve the best performance, with Random Forest (L4) achieving the best cost reduction of 70.8%. The Bayesian Structural Causal Model (L7) achieved competitive financial performance of a 66.4% cost reduction. On the other hand, it is capable of attributing predicted failures to specific types without requiring failure-type labels during training. This capability can prove to be a favorable trade-off in predictive maintenance scenarios.
@@
    Ensemble methods as performance ceiling: Tree-based ensemble models (L4, L5, L6) consistently outperform single models including the decision tree that mirrors the true data-generating process. This demonstrates the strength of ensembling methods in reducing variance and preventing overfitting.
@@
    Causal AI for interpretability: The Bayesian SCM (L7) delivers competitive cost savings and provides actionable insights in the form of failure attribution. The performance gap relative to the best ensemble model is modest and may be outweighed by the practical value of the failure diagnosis.
@@
    Failure attribution without attribution labels: The Bayesian SCM (L7) correctly attributes predicted failures to specific physical mechanisms (TWF, HDF, PWF, OSF) despite being trained only on binary failure labels. This means there is no additional labeling cost required for training the model.
@@
    Robust generalization: The Bayesian SCM demonstrates solid generalization (3.17 percentage point train-test gap), which is consistent with the expectation that causal relationships are invariant properties of the physical system and less susceptible to distributional variation in training data.
@@
    Static analysis: This study evaluates models on synthetic data without real-world deployment. Practical implementation would require monitoring for distribution drift, model degradation, and integration with existing maintenance management systems.
@@
Future research directions include: (1) validation on operational data; (2) investigation of causal discovery algorithms to automate DAG construction; (3) development of online learning approaches for continual model adaptation; (4) integration with IoT sensor networks and SCADA systems for real-time deployment; and (5) extension to multi-component systems with cascading failure modes.
6. Author Contributions
The authors are listed in alphabetical order.
Krishna Taduri performed exploratory data analysis, developed the causal structural model (including failure attribution) and integrated cost optimization.
Shaunak Dhande performed exploratory data analysis and developed the correlation-based models.
Chutian Ma reviewed and revised the manuscript, and contributed to the discussion of the failure attribution mechanism.
Giacinto Paolo Saggese and Paul Smith provided overall conceptual guidance, contributed to architectural design decisions, and offered feedback on the manuscript.